OpenStudy (anonymous):

The perimeter of an equilateral triangle is 26 after each of the sides is increased by 2 units. If S is the length of the side of the original triangle, write an equation for the new perimeter in terms of S.

OpenStudy (anonymous):

Perimeter of equilateral triangle = 3a 26 = 3(S+2) 26 = 3S + 6 20 = 3S (area of original triangle)

OpenStudy (triciaal):